OpenStudy (help):

For each of the following angles, find the radian mesaure of hte angle with the given degree measure: -150º -130º -20º 360º

OpenStudy (anonymous):

use the proportion pi/180 = n/(any degree) so\[\pi/180 = n/150;then; 150\pi/180=n\] = 5pi/6, the rest are set up all the same way.

OpenStudy (help):

Thanks a million!

OpenStudy (anonymous):

no sweat cheif, keep your head in there.